OpenStudy (anonymous):

x^2+10x+16/x^2+6x+8 divided by 1/x+4

OpenStudy (anonymous):

its 2x +8

OpenStudy (anonymous):

I got (x+8)(x+2)/(x+6)(x+2) x x+4/1

OpenStudy (anonymous):

This, right? \[ \frac{\frac{x^2+10x+16}{x^2+6x+8}}{\frac{1}{x+4}} \] Dividing fractions is like multiplying by the reciprocal, so that above equals: \[ \frac{(x+4)(x^2+10x+16)}{x^2+6x+8} \] The denominator factors into: \[ x^2 + 6x + 8 = (x+4)(x+2) \] Also, the numerator factors into \[ x^2 + 10x + 16 = (x+8)(x+2) \] So now, the (x+4)s and the (x+2)s cancel from both the numerator and the denominator, and the final answer is just x+8.